# EMAC+: Embodied Multimodal Agent for Collaborative Planning with VLM+LLM

### v1

@@ 4.1 Experiments Setup
We base our experiments on two embodied environments, ALFWorld (Shridhar2020ALFWorldAT, ) and RT-1 (Brohan2022RT1RT, ). The former environment provides a cross-modality simulation platform that encompasses a wide range of embodied household tasks. RT-1 has three robot environments, including a Task and Motion Planning (TAMP) domain where a robot has to manipulate (grasp and stack) objects, a tabletop pushing environment, and a mobile manipulation domain. We provide more explanations and details for each environment as follows:
@@
ALFWorld. For each task in ALFWorld, the simulator integrates a visual environment, rendered by the AI2THor (Kolve2017AI2THORAI, ), with a corresponding textual environment. The textual environment employs PDDL (Helmert2006TheFD, ) to translate each pixel observation from the simulator into an equivalent text-based observation, and then construct an interactive world using the TextWorld engine (Ct2018TextWorldAL, ). The tasks within the benchmark are categorized into six types: “Pick & Place”, “Clean & Place”, “Heat & Place”, “Cool & Place”, “Look in Light, and Pick Two Objects & Place”. Each task requires an agent to execute a series of text-based actions, such as “go to safe 1”, “open safe 1”, or “heat egg 1 with microwave 1”, following a predefined instruction. These actions involve navigating and interacting with the environment. A human expert might need more than 30 steps to solve a job in this benchmark that involves interacting with more than 10 objects. This makes it hard for an agent to plan ahead, comply with instructions, and use common sense knowledge. To make the comparison fair, we use the same setting as earlier work and use 134 out-of-distribution (OOD) jobs to test all of the baselines.
@@
RT-1. RT-1 has three domains. We train our MMA on expert data from each domain of RT-1. In many cases, this is a sparse amount of data per task. The TAMP problems require a lot of combinatorial analysis of possible plans, and many decision sequences can’t be made. MMA must formulate strategies comprising several stages, characterized by intricate decision limits. The multi-object tabletop pushing environment is derived from the publicly accessible Language-Table dataset (Lynch2022InteractiveLT, ) and presents challenges due to the presence of several objects, a high cardinality of language, and intricate pushing dynamics. In both the TAMP and Language-Table environments, our MMA must analyze the positions of the objects. Merely identifying the things on the table and their approximate relationships is inadequate; a more nuanced understanding of the scene’s geometry is essential for task resolution.
@@
Baselines. To verify the effectiveness of the collaborative learning scheme, we compare our MMA with several baselines and SOTA agents. We categorize the baselines with three categories: vision models, language models (LLM agents), and vision-language models (VLM agents). For vision models, we pick ResNet-18 He2015DeepRL and MCNN-FPN He2017MaskR , which utilize pretrained encoders to extract salient features from each pixel observation. Then, behavior cloning is used to train a Multi-Layer Perceptron (MLP) policy on a demonstration dataset that was already made using the extracted features as background. In contrast to vision models that operate inside a visual environment, language models do analogous tasks in a parallel, text-based framework. BUTLER (Shridhar2020ALFWorldAT, ) utilizes a transformer sequence-to-sequence model augmented with a pointer softmax method (Glehre2016PointingTU, ). This design compiles prior observations as input to produce text-based actions incrementally, one token at a time. GPT-BUTLER (Micheli2021LanguageMA, ), a variation of the GPT-2 model (Radford2019LanguageMA, ), is initially pretrained on a static demonstration dataset and subsequently finetuned with data gathered online. ReAct (Yao2022ReActSR, ) employs an innovative method by leveraging LLMs to produce reasoning traces and task-specific actions concurrently. This strategy assists the agent in formulating, monitoring, and revising its action plans interactively. Reflexion (Shinn2023ReflexionLA, ) similarly utilizes a large language model, concentrating on the analysis of environment feedback. It retains this reflective text in an episodic memory buffer, augmenting the agent’s capacity to refine actions in future trials. DEPS (Wang2023DescribeEP, ) rectifies inaccuracies in prior LLM-generated actions by incorporating descriptions of the action execution process and offering self-explanations for the feedback. Furthermore, beyond the single-agent paradigm, AutoGen (Wu2023AutoGenEN, ) demonstrates the capability to do a wide array of tasks through the collaboration of numerous LLM agents. We examine a variety of vision-language models, including MiniGPT-4 (Zhu2023MiniGPT4EV, ), BLIP-2 (Li2023BLIP2BL, ), LLaMA-Adaptor (Gao2023LLaMAAdapterVP, ), and InstructBLIP (Dai2023InstructBLIPTG, ), as agents for engaging with the visual environment. In contrast to pure vision or language models, VLMs are engineered to interpret and synthesize both visual and verbal information, providing a more comprehensive understanding of the environment. We finetune these agents on a pre-collected demonstration dataset to align them with the unique needs of the ALFWorld benchmark. In RT-1, We pick PaLM-E (Driess2023PaLMEAE, ), SayCan (Ahn2022DoAI, ), and PaLI (Chen2022PaLIAJ, ) as baselines, since the RT-1 environment requires the agent to generate low-level control actions rather than textual actions.
@@ 4.2 ALFWorld Environment
The architectural design of our MMA is shown in Fig. 1. The core of the VLM part in MMA employs a Query Transformer (Q-former) to process the visual data. a. This Q-Former extracts features using a frozen ViT encoder. The output of this Q-Former comprises 32 visual tokens, which undergo a linear projection layer before reaching an LLM decoder. In our experiments, we choose text-davinci-003 developed by OpenAI as the LLM expert due to its outstanding performance in reasoning and planning tasks (Shinn2023ReflexionLA, ; Wang2023DescribeEP, ; Wu2023AutoGenEN, ; Yao2022ReActSR, ).
@@
We find the EMAC+ outperforms all baselines on the ALFWorld benchmark with a lower training cost. We compare EMAC+ with 13 other representative agents on both textual and visual environments and report the results in Table. 1. We evaluate two main metrics: the success rate, which is the percentage of successful trials, and the average number of interaction steps needed to finish the task. A lower number of steps means the task was completed more quickly. EMAC+ has exceptional performance across both criteria, surpassing all VLM agents in visual contexts. Also, EMAC+’s performance is much better than that of VM agents, which shows how important it is for VLMs to have prior information built in. Interestingly, EMAC+’s performance is analogous to that of LLM agents utilizing precise semantic representations of visual observations. We attribute this to the training strategy, which involves imitating from an LLM expert agent instead of learning from scratch in a purely visual setting. Consequently, EMAC+ stands out as a VLM agent that substantially surpasses SOTA LLM agents, such as AutoGen (Wu2023AutoGenEN, ) and DEPS (Wang2023DescribeEP, ).
@@ 4.3 RT-1 Environment
For RT-1 robot tasks, we run our experiments in a simulator on the RT-1 codebase (Brohan2022RT1RT, ). We report the results on motion planning tasks, table pushing tasks and manipulation tasks in Table. 2 - Table. 3.
@@
Table 2: Results on RT-1 TAMP environment. where data from TAMP constitutes only $1\%$ (i.e., 320 samples for $p_{1},p_{2}$ each) of total training data size. PaLM-E (OSRT) despites using no large-scale data, provides the most effective input encodings for learning. PaLM-E (states) provides ground-truth object-centric information during training. The LLM agent in all PaLM-E baselines is frozen (following the original settings in  Driess2023PaLMEAE ).
RT-1 TAMP environment. Table. 2 reports the success rates and VQA performance in the TAMP environment. Note that the LLM in baseline methods is frozen (for pre-trained LLM). For the pre-training LLM, we use the dataset collected in  (Padalkar2023OpenXR, ), which contains 90000$\sim$ training scenes in the TAMP environment. The results in Table. 2 are tested when trained on $1\%$ of the dataset, corresponding to 300 examples for the two planning tasks. This setting is aligned with it in  Driess2023PaLMEAE . Most input representations exhibit comparable performance for 3-5 objects in the scene, equivalent to the quantity in the training set.  Driess2023PaLMEAE discussed the advantages of pre-training ViT with fully mixed data (text + visual), which will bring significant improvement. EMAC+ follows this pretraining approach. We found that although there is not much difference between EMAC+ and PaLM-E in visual QA, EMAC+ far outperforms PaLM-E in planning. This can be ascribed to EMAC+’s capacity to synchronize textual and visual states. In other words, both have employed pixel and textual observations during pretraining. During the training phase, EMAC+ can get deeper insights into environment dynamics by observing the execution of actions by the VLM agent and the LLM agent. This information is irreplaceable for action planning.
@@
RT-1 Planning tasks. Table. 2 reports the performance on long-horizon tasks from the Language-table environment (Lynch2022InteractiveLT, ). Refer to Table. 4 for the task prompts in each planning task. The results are evaluated under a few-shot regime (10 demos per task) to assess the agent’s generalization capability in out-of-distribution (OOD) tasks. Our findings indicate that EMAC+, possessing merely 7B parameters (Vicuna-7B), significantly surpasses PaLM-12B. The collaborative structure of EMAC+ significantly enhances the generalization capability of the LLM agent for unseen tasks, whereas SyaCan and PaLI are ineffective and incapable of addressing any test tasks.
@@
Evaluation metrics for manipulation tasks (Table. 3). In RT-1 manipulation tasks, we mainly follow the setup in  Ahn2022DoAI , where the robot needs to plan a sequence of navigation and manipulation actions based on the task instruction and pixel observations. We have two test metrics to evaluate the performance, i.e., affordance prediction and failure detection. Affordance prediction is designed to investigate EMAC+’s performance on predicting whether a skill of the low-level policy can be executed in the environment. More specifically, given an <image> and a question “Does the control policy can do <skill> here?”. Failure detection is commonly used for robotic closed-loop planning to detect failures. For example, given an <image>, the model needs to understand “Is <skill> successfully executed by the robot?”. This is an important attribute of the multi-modal agent. It is theoretically essential for the agent to comprehend the reasons for the failure of the current task. It provides guarantee for the LLM’s ability to modify plans dynamically.
4.4 Ablation Studies
[…]
right corner}. Push that block to the other block of
the same color.
[…]
corners?
[…]
the {left/right} side together, without bringing over
any of the blocks that are on the {right/left} side?
Figure 4: Task prompts for planning tasks. (Table. 2)
@@ 5 Conclusion and limitations
Appendix A Prompts for LLM expert in EMAC+
A.1 Translated Textual state and LLM prompts in ALFWorld
We supply all LLM prompts for the training process (Alg. 1) of EMAC+. We utilize the prompting strategy established by ReAct [42] while disregarding the reasoning traces, specifically the "think" stages, throughout the imitation learning process between EMAC+ and the LLM actor. Following each trial $i$, the retrospective feedback $\mathcal{P}_{i}$ produced by the LLM critic will be incorporated into the long-term memory $\mathcal{P}$. In practice, we limit $\mathcal{P}$ to a maximum of 1-3 stored feedback instances to comply with the maximum context duration of the LLM. The textual state and task instruction is shown in Table. 3. The LLM plans the action sequence to complete the task as shown in Table. 4. The agent executes the planned actions in the environment, then the LLM generates the retrospective feedback based on the environment feedback (Table. 5). The LLM expert will re-plann actions if the agent cannot successfully execute the previous actions (Table. 6).
[…]
@@ A.1 Translated Textual state and LLM prompts in ALFWorld
[…]
@@
[…]
@@
[…]
@@
long-term memory $\mathcal{P}:$ [$\cdots$]
@@
[…]
@@
long-term memory $\mathcal{P}:$ [$\cdots$]
@@ Appendix C The structure of EMAC+
We list the structure of EMAC+ and hyperparameters in Table.9 and Table. 10. These hyperparameters are primarily based on those suggested for fine-tuning the InstructBLIP model [7]. During training, we exclusively change the parameters of the linear projection layer while maintaining the remaining components in a frozen state, consistent with the approach utilized in instruction tuning for numerous existing studies [12, 45]. We employ the AdamW optimizer, initiating with a linear warmup of the learning rate, succeeded by a linear decay, culminating in a minimal learning rate of 0. Furthermore, we eliminate the instruction input of Q-Former, utilized in InstructBLIP, and observe an enhancement in performance across all studies.
@@
LLM decoder Vicuna-7b-v1.1 [43]
Image encoder ViT-L  [31]
Q-Former BERT base ${}_{\text{base }}$ [9]
Pretrained weights InstructBLIP  [7]
@@
Base model for LLM expert text-davinci-003
@@
LLM decoder Vicuna-7b-v1.1 [43]
Image encoder ViT-L  [31]
Q-Former BERT base ${}_{\text{base }}$ [9]
Pretrained weights InstructBLIP  [7]
@@
Base model for LLM expert text-davinci-003

### v2

@@ 4.1 Experiments Setup
We base our experiments on two embodied environments, ALFWorld (Shridhar et al., 2020) and RT-1 (Brohan et al., 2022). The former environment provides a cross-modality simulation platform that encompasses a wide range of embodied household tasks. RT-1 has three robot environments, including a Task and Motion Planning (TAMP) domain where a robot has to manipulate (grasp and stack) objects, a tabletop pushing environment, and a mobile manipulation domain. We provide more explanations and details for each environment as follows:
@@
ALFWorld. For each task in ALFWorld, the simulator integrates a visual environment, rendered by the AI2THor (Kolve et al., 2017), with a corresponding textual environment. The textual environment employs PDDL (Helmert, 2006) to translate each pixel observation from the simulator into an equivalent text-based observation, and then construct an interactive world using the TextWorld engine (Côté et al., 2018). The tasks within the benchmark are categorized into six types: “Pick & Place”, “Clean & Place”, “Heat & Place”, “Cool & Place”, “Look in Light, and Pick Two Objects & Place”. Each task requires an agent to execute a series of text-based actions, such as “go to safe 1”, “open safe 1”, or “heat egg 1 with microwave 1”, following a predefined instruction. These actions involve navigating and interacting with the environment. A human expert might need more than 30 steps to solve a job in this benchmark that involves interacting with more than 10 objects. This makes it hard for an agent to plan ahead, comply with instructions, and use common sense knowledge. To make the comparison fair, we use the same setting as earlier work and use 134 out-of-distribution (OOD) jobs to test all of the baselines.
@@
RT-1. RT-1 has three domains. We train EMAC+ on expert data from each domain of RT-1. In many cases, this is a sparse amount of data per task. The TAMP problems require a lot of combinatorial analysis of possible plans, and many decision sequences can’t be made. EMAC+ must formulate strategies comprising several stages, characterized by intricate decision limits. The multi-object tabletop pushing environment is derived from the publicly accessible Language-Table dataset (Lynch et al., 2022) and presents challenges due to the presence of several objects, a high cardinality of language, and intricate pushing dynamics. In both the TAMP and Language-Table environments, EMAC+ must analyze the positions of the objects. Merely identifying the things on the table and their approximate relationships is inadequate; a more nuanced understanding of the scene’s geometry is essential for task resolution.
@@
Baselines. To verify the effectiveness of the collaborative learning scheme, we compare EMAC+ with several baselines and SOTA agents. We categorize the baselines with three categories: vision models, language models (LLM agents), and vision-language models (VLM agents). For vision models, we pick ResNet-18 He et al. (2015) and MCNN-FPN He et al. (2017), which utilize pretrained encoders to extract salient features from each pixel observation. Then, behavior cloning is used to train a Multi-Layer Perceptron (MLP) policy on a demonstration dataset that was already made using the extracted features as background. In contrast to vision models that operate inside a visual environment, language models do analogous tasks in a parallel, text-based framework. Qwen (Bai et al., 2023) achieves competitive performance on diverse benchmarks and illustrates its strong potential on general tasks. BUTLER (Shridhar et al., 2020) utilizes a transformer sequence-to-sequence model augmented with a pointer softmax method (Çaglar Gülçehre et al., 2016). This design compiles prior observations as input to produce text-based actions incrementally, one token at a time. GPT-BUTLER (Micheli & Fleuret, 2021), a variation of the GPT-2 model (Radford et al., 2019), is initially pretrained on a static demonstration dataset and subsequently finetuned with data gathered online. ReAct (Yao et al., 2022) employs an innovative method by leveraging LLMs to produce reasoning traces and task-specific actions concurrently. This strategy assists the agent in formulating, monitoring, and revising its action plans interactively. Reflexion (Shinn et al., 2023) similarly utilizes a large language model, concentrating on the analysis of environment feedback. It retains this reflective text in an episodic memory buffer, augmenting the agent’s capacity to refine actions in future trials. DEPS (Wang et al., 2023) rectifies inaccuracies in prior LLM-generated actions by incorporating descriptions of the action execution process and offering self-explanations for the feedback. Furthermore, beyond the single-agent paradigm, WALL-E (Zhou et al., 2024) has SOTA performance on ALFWorld across various LLM-based agents, and AutoGen (Wu et al., 2023) demonstrates the capability to do a wide array of tasks through the collaboration of numerous LLM agents. We examine a variety of vision-language models, including MiniGPT-4 (Zhu et al., 2023), BLIP-2 (Li et al., 2023b), LLaMA-Adaptor (Gao et al., 2023b), and InstructBLIP (Dai et al., 2023), as agents for engaging with the visual environment. In contrast to pure vision or language models, VLMs are engineered to interpret and synthesize both visual and verbal information, providing a more comprehensive understanding of the environment. We finetune these agents on a pre-collected demonstration dataset to align them with the unique needs of the ALFWorld benchmark. In RT-1, We pick PaLM-E (Driess et al., 2023), SayCan (Ahn et al., 2022), and PaLI (Chen et al., 2022) as baselines, since the RT-1 environment requires the agent to generate low-level control actions rather than textual actions.
@@ 4.2 ALFWorld Environment
The architectural design of EMAC+ is shown in Fig. 1. The core of the VLM part in EMAC+ employs a Query Transformer (Q-former) to process the visual data. a. This Q-Former extracts features using a frozen ViT encoder. The output of this Q-Former comprises 32 visual tokens, which undergo a linear projection layer before reaching an LLM decoder. In our experiments, we choose LLaMa3 as the LLM expert due to its outstanding performance in reasoning and planning tasks (Shinn et al., 2023; Wang et al., 2023; Wu et al., 2023; Yao et al., 2022).
@@
AdaPlanner 0.91(-) 1.00(-) 1.00(-) 0.89(-) 1.00(-) 0.97(-) 0.47(-)
Qwen2-Chat 0.67(-) 0.62 (-) 0.75 (-) 0.68(-) 0.61(-) 0.78 (-) 0.52 (-)
WALL-E 0.95(-) 1.00(-) 0.97(-) 1.00(-) 0.86(-) 0.83(-) 0.95(-)
VLMs Agent
@@
We find the EMAC+ outperforms all baselines on the ALFWorld benchmark with a lower training cost. We compare EMAC+ with 13 other representative agents on both textual and visual environments and report the results in Table. 1. We evaluate two main metrics: the success rate, which is the percentage of successful trials, and the average number of interaction steps needed to finish the task. A lower number of steps means the task was completed more quickly. EMAC+ has exceptional performance across both criteria, surpassing all VLM agents in visual contexts. Additionally, EMAC’s performance is significantly better than that of VM agents, highlighting the importance of having prior information built in for VLMs. Interestingly, EMAC+’s performance is analogous to that of LLM agents utilizing precise semantic representations of visual observations. We attribute this to the training strategy, which involves imitating from an LLM expert agent instead of learning from scratch in a purely visual setting. Consequently, EMAC+ stands out as a VLM agent that substantially surpasses SOTA LLM agents, such as AutoGen (Wu et al., 2023) and DEPS (Wang et al., 2023).
@@ 4.3 RT-1 Environment
For RT-1 robot tasks, we run our experiments in a simulator on the RT-1 codebase (Brohan et al., 2022). We report the results on motion planning tasks, table pushing tasks and manipulation tasks in Figure. 2 - Figure. 3.
[…]
@@
Table 2: Results on RT-1 TAMP environment. where data from TAMP constitutes only $1\%$ (i.e., 320 samples for $p_{1},p_{2}$ each) of total training data size. PaLM-E (OSRT) despites using no large-scale data, provides the most effective input encodings for learning. PaLM-E (states) provides ground-truth object-centric information during training. The LLM agent in all PaLM-E baselines is frozen (following the settings in  Driess et al. (2023)).
@@
RT-1 Planning tasks. Figure. 2 reports the performance on long-horizon tasks from the Language-table environment (Lynch et al., 2022). Refer to Table. 5 for the task prompts in each planning task. The results are evaluated under a few-shot regime (10 demos per task) to assess the agent’s generalization capability in out-of-distribution (OOD) tasks. Our findings indicate that EMAC+, possessing merely 7B parameters (Vicuna-7B), significantly surpasses PaLM-12B. The collaborative structure of EMAC+ significantly enhances the generalization capability of the LLM agent for unseen tasks, whereas SyaCan and PaLI are ineffective and incapable of addressing any test tasks.
4.4 Ablation Studies
@@ 5 Conclusion and limitations
Ethics Statement
This work investigates embodied multimodal agents in simulated environments (ALFWorld and RT-1) and does not involve human subjects, personal data, or real-world robotic deployments. Therefore, there are no direct ethical risks in the experiments presented. We acknowledge, however, that embodied AI systems deployed in real-world settings could introduce safety risks, particularly if planning or perception errors occur. To mitigate such risks, we emphasize that EMAC+ is a research prototype evaluated in controlled simulations. We encourage future work to incorporate safety constraints, human oversight, and rigorous evaluation before applying such systems in physical environments.
Appendix A Experiments Details for EMAC+
A.1 Evaluation metrics for manipulation tasks (Table. 3).
In RT-1 manipulation tasks, we mainly follow the setup in  Ahn et al. (2022), where the robot needs to plan a sequence of navigation and manipulation actions based on the task instruction and pixel observations. We have two test metrics to evaluate the performance, i.e., affordance prediction and failure detection. Affordance prediction is designed to investigate EMAC+’s performance on predicting whether a skill of the low-level policy can be executed in the environment. More specifically, given an <image> and a question “Does the control policy can do <skill> here?”. Failure detection is commonly used for robotic closed-loop planning to detect failures. For example, given an <image>, the model needs to understand “Is <skill> successfully executed by the robot?”. This is an important attribute of the multi-modal agent. It is theoretically essential for the agent to comprehend the reasons for the failure of the current task. It provides guarantee for the LLM’s ability to modify plans dynamically.
Task 1: Q1: There is a block that is closest to {i.e., top right corner}. Push that block to the other block of the same color.
Task 2: Q2: How to sort the blocks by colors into corners?
Task 3: Q3: How to push all the blocks that are on the {left/right} side together, without bringing over any of the blocks that are on the {right/left} side?
Figure 5: Task prompts for planning tasks. (Figure. 2)
A.2 Translated Textual state and LLM prompts in ALFWorld
We supply all LLM prompts for the training process (Alg. 1) of EMAC+. We utilize the prompting strategy established by ReAct (Yao et al., 2022) while disregarding the reasoning traces, specifically the "think" stages, throughout the imitation learning process between EMAC+ and the LLM actor. Following each trial $i$, the retrospective feedback $\mathcal{P}_{i}$ produced by the LLM critic will be incorporated into the long-term memory $\mathcal{P}$. In practice, we limit $\mathcal{P}$ to a maximum of 1-3 stored feedback instances to comply with the maximum context duration of the LLM. The textual state and task instruction is shown in Table. 3. The LLM plans the action sequence to complete the task as shown in Table. 4. The agent executes the planned actions in the environment, then the LLM generates the retrospective feedback based on the environment feedback (Table. 5). The LLM expert will re-plann actions if the agent cannot successfully execute the previous actions (Table. 6).
Environment: You are in the middle of a room. Looking quickly around you, you see a cabinet 4, a cabinet 3, a cabinet 2, a cabinet 1, a countertop 1, a garbagecan 1, a handtowelholder 2, a handtowelholder 1, a sinkbasin 2, a sinkbasin 1, a toilet 1, a toiletpaperhanger 1, and a towelholder 1.
@@ A.2 Translated Textual state and LLM prompts in ALFWorld
> think: To solve the task, I need to find and take a sparybottle, then put it on toilet.
> think: First I need to find a spraybottle. A spraybottle is more likely to appear in cabinet (1-4), countertop (1), toilet (1), sinkbasin (1-2), garbagecan (1). I can check one by one, starting withcabinet 1.
@@
Env. feedback: On the cabinet 1, there is cloth 1, a soapbar 1, a soapbottle 1
> think: On the cabinet 1, I can see cloth 1, a soapbar 1, a soapbottle 1. I can continue to execute step 2.
@@
Env. feedback: cabinet 2
> think: cabinet 2 is closed. I can continue to execte step 3.
@@
long-term memory $\mathcal{P}:$ $\cdots$]
@@
> think: step 3 is failed to execute. cabinet 2 is closed.
@@
long-term memory $\mathcal{P}:$ $\cdots$]
@@ Appendix C The structure of EMAC+
We list the structure of EMAC+ and hyperparameters in Table.9 and Table. 10. These hyperparameters are primarily based on those suggested for fine-tuning the InstructBLIP model (Dai et al., 2023). During training, we exclusively change the parameters of the linear projection layer while maintaining the remaining components in a frozen state, consistent with the approach utilized in instruction tuning for numerous existing studies (Gao et al., 2023b; Zhu et al., 2023). We employ the AdamW optimizer, initiating with a linear warmup of the learning rate, succeeded by a linear decay, culminating in a minimal learning rate of 0. Furthermore, we eliminate the instruction input of Q-Former, utilized in InstructBLIP, and observe an enhancement in performance across all studies.
@@
LLM decoder Vicuna-7b-v1.1 Zheng et al. (2023)
Image encoder ViT-L  Radford et al. (2021)
Q-Former BERT $_{\text{base }}$ Devlin et al. (2019)
Pretrained weights InstructBLIP  Dai et al. (2023)
@@
Base model for LLM expert LLaMa3
@@
LLM decoder Vicuna-7b-v1.1 Zheng et al. (2023)
Image encoder ViT-L  Radford et al. (2021)
Q-Former BERT $_{\text{base }}$ Devlin et al. (2019)
Pretrained weights InstructBLIP  Dai et al. (2023)
@@
Base model for LLM expert LLaMa3